OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x2 + 3x + 2 y = x2 + 2x + 1 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

OpenStudy (anonymous):

y = x^2 +3x + 2 y = x^2 +2x +x +2 y = (x+2)(x+1) 0 = x+2, x= -2 0 = x+1, x= -1 X intercepts are at (-2, 0) and (-1, 0) y = x^2 + 2x + 1 y = x^2 +x +x +1 y = (x+1)(x+1) or y = (x+1)^2 0 = x+1, x= -1 X intercept is at (-1,0) Both equations represent parabolas, and both cross the X axis at (-1,0). The first equation has possible negative x values whereas the second does not.

OpenStudy (anonymous):

thts 1st part

OpenStudy (anonymous):

oh hahaha thankya

OpenStudy (anonymous):

In addition to that, they have different factors. Similarities 1. Both quadratics 2. Both intersect at same point (-1,0) Differences: 1. Different axes of symmetry 2. Different x and y intercept 3. Different factors

OpenStudy (anonymous):

thts it

OpenStudy (anonymous):

the writing from the 1st part i gave u and this also is part 2

OpenStudy (anonymous):

and part 1

OpenStudy (anonymous):

May I help ?:)

OpenStudy (anonymous):

Solve 3x2 = 12x. x = 0, x = 4 x = 0, x = −4 x = 3, x = 4 x = 3, x = −4

OpenStudy (anonymous):

wait... let me check part 2 fully

OpenStudy (anonymous):

ok and this is aprt 2: The two equations share the following common properties: They both have a root at x = -1. They also both open upwards. The difference between the two equations is that the second equation is a perfect square which is tangent to the x-axis at (-1,0) and does not take on any negative values. The equations both have distinct turning points and axis of symmetry.

OpenStudy (anonymous):

x = 2, x= -2

OpenStudy (anonymous):

o 12xz

OpenStudy (anonymous):

x = 0, x=4

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

3x^2-12x+0=0 So we have Quadratic equation . So use from Delta and find x1 and x2 .

OpenStudy (anonymous):

What are the solutions of 6x2 + 7x – 5 = 0? solve for the x's

OpenStudy (anonymous):

x = negative 5/3, x = 1/2

OpenStudy (anonymous):

Identify which equation is a quadratic equation. y = −5x − 3 y = −4x(x + 10) + 34 y + x2 = (x − 4)(x + 3) y − 2 = (x − 2)(x2 + 1)

OpenStudy (anonymous):

y = −4x(x + 10) + 34

OpenStudy (anonymous):

wait let me double check

OpenStudy (anonymous):

@samsterz : i think say the formula is better .

OpenStudy (anonymous):

yes its y = −4x(x − 4) + 2x

OpenStudy (anonymous):

Determine whether the graph of y = 3x2 − 7x + 2 opens up or down and whether it has a maximum or minimum point. Down and Maximum Down and Minimum Up and Maximum Up and Minimum

OpenStudy (anonymous):

wait lolol wrong math

OpenStudy (anonymous):

@jordynnn : Do u know that formula ???

OpenStudy (anonymous):

no, i suck at math :(

OpenStudy (anonymous):

jordynn did u submit a quiz or something, if so what did u get prior?>

OpenStudy (anonymous):

answer is : Up; Minimum

OpenStudy (anonymous):

i havent submitted anything yet

OpenStudy (anonymous):

lol it must be long

OpenStudy (anonymous):

When the quadratic equation y = (x − 2)(3x + 1) is written in standard form, what is the value of the coefficient "b"? −5 3 −2 5

OpenStudy (anonymous):

omg it is

OpenStudy (anonymous):

answer is A: -5

OpenStudy (anonymous):

What is the vertex for the graph of y = −2x2 + 8x − 13? (2, 11) (2, −5) (−2, −37) (−2, −21)

OpenStudy (anonymous):

(2, −5)

OpenStudy (anonymous):

What are the solutions of x2 − 3x − 18 = 0? x = 3, x = −6 x = −3, x = 6 x = 3, x = 6 x = −3, x = −6

OpenStudy (anonymous):

@samsterz : What are you doing ??????

OpenStudy (anonymous):

?? what do u mean

OpenStudy (anonymous):

is it C?

OpenStudy (anonymous):

You just answer to her . She can learn them ...

OpenStudy (anonymous):

x= -3, x=6

OpenStudy (anonymous):

lololol

OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − x2 = 11x + 30? (−5, 0) and (6, 0) (5, 0) and (−6, 0) (−5, 0) and (−6, 0) (5, 0) and (6, 0)

OpenStudy (anonymous):

its greatly appreciated that @samsterz is helping me

OpenStudy (anonymous):

x intercepts (-6,0),(-5,0)

OpenStudy (anonymous):

im on the last one now!! but its long again .. ;/

OpenStudy (anonymous):

lol show it

OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (anonymous):

1st part : x2-25=0 (x+5)(x-5)=0 x=-5 or x=5 x2=2x+15 x2-2x-15=0 (x-5)(x+3)=0 x=5 or x=-3 part 2 and 3: 2. The values of 'x' are the points on the x-axis where the curve intercepts(crosses) . 3. They are both 'bowl' shaped. They both have the same coefficient of 'x^2' so they both have the same degree of 'closure'. The first has a minimum at x = 0 and y = -5 The second has a minimum at x = 1 and y = -16 . It also intercepts the y-axis at y = -15. This graph opens down. I know because it has a negative number with the x^2 term. The graph has a maximum, as do all graphs that open down. The vertex is at (0.7, 0.45) The x-intercepts are at (0.4, 0) and (1, 0)

OpenStudy (anonymous):

there was a second one to it that i forgot

OpenStudy (anonymous):

Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following: Does the graph open up or down? How do you know? (2 points) Explain whether the graph has a maximum or minimum point. (2 points) Find the vertex and x-intercepts of the graph. (3 points)

OpenStudy (anonymous):

Here's one for you: y = x^2 - 8x + 12 This factors as y = (x - 6)(x - 2) So, x-intercepts are at (6, 0), and (2, 0) This opens up because the implied "a" value is 1. It also has the vertex form of y = (x - 4)^2 - 4 so it has a vertex at (4, -4) and it is at this point where the graph and equation has its low-point. done.

OpenStudy (anonymous):

lol :P

OpenStudy (anonymous):

i got a 92% :D

OpenStudy (anonymous):

;(

OpenStudy (anonymous):

i missed like 2?

OpenStudy (anonymous):

no, i did some before asking you, you got them ALL

OpenStudy (anonymous):

hopefully ur score will go up when u the teacher grades written

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

you're amazing thankyou so much! :)

OpenStudy (anonymous):

yup, trying to do some good deeds